if velocity v of a particle moving along a straight line decreases linearly with its position coordinates s from 50m/sec to zero at s=100mters,find the time taken to 100 metrs?

Grade:12

1 Answers

Rohan Das
10 years ago

here,u=50m/s, v=0m/s, s=100m

we know v*v-u*u=2as

or a=(v*v- u*u)/2s

or a=25ms-2

also v=u+at

or t=(v-u)/a

or t=2s